# Visual Enhanced Depth Scaling for Multimodal Latent Reasoning

### 16 Apr 2026 v2

@@ Abstract
Multimodal latent reasoning has emerged as a promising paradigm that replaces explicit Chain-of-Thought (CoT) decoding with implicit feature propagation, simultaneously enhancing representation informativeness and reducing inference latency. By analyzing token-level gradient dynamics during latent training, we reveal two critical observations: (1) visual tokens exhibit significantly higher and more volatile gradient norms than their textual counterparts due to inherent language bias, resulting in systematic visual under-optimization; and (2) semantically simple tokens converge rapidly, whereas complex tokens exhibit persistent gradient instability constrained by fixed architectural depths. To address these limitations, we propose a visual replay module and routing depth scaling to collaboratively enhance visual perception and refine complicated latents for deeper contextual reasoning. The former module leverages causal self-attention to estimate token saliency, reinforcing fine-grained grounding through spatially-coherent constraints. Complementarily, the latter mechanism adaptively allocates additional reasoning steps to complex tokens, enabling deeper contextual refinement. Guided by a curriculum strategy that progressively internalizes explicit CoT into compact latent representations, our framework achieves state-of-the-art performance across diverse benchmarks while delivering substantial inference speedups over explicit CoT baselines.
@@ 1 Introduction
Figure 1: Panel (a) depicts the token-wise Frobenius norm of gradients within different layer throughout the overall training process. Notably, visual tokens (i.e., those with indices within about $\left[200,280\right]$) exhibit consistently larger gradient magnitudes accompanied by abrupt spikes, suggesting that they are more challenging to optimize compared to textual tokens (indices without $\left[200,280\right]$). Panels (b) and (c) further reveal distinct evolution patterns of the Nuclear norm in the QKVO projection matrix across layers and training epochs: gradients for easy samples decay rapidly and converge smoothly, whereas hard samples maintain elevated gradient norms with persistent oscillations. Here, $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$ denote the low-rank matrix in LoRA [hu2022lora].
Over the past few years, large language models (LLMs) have achieved remarkable progress in complex reasoning, propelled by scaling laws in data volume and model capacity [DBLP:journals/corr/abs-2001-08361]. Advanced techniques such as Chain-of-Thought (CoT) prompting [zhang2023multimodal, shao2024visual] and reinforcement learning (RL) [deepseekr1] for trajectory optimization have proven highly effective in text-only domains. Extending these capabilities to the multimodal realm has thus become a pivotal research direction. Current approaches primarily follow three paradigms. First, Text-based Reasoning [zhang2023multimodal, MitraCCoT, lei2024scaffolding] generates explicit multi-step textual chains before producing an answer. However, these methods typically rely on static visual inputs, and recent studies [lookback2025, DBLP:journals/corr/abs-2502-17425] indicate that visual grounding deteriorates significantly over extended reasoning chains. Second, Tool-augmented Reasoning manipulates visual inputs through external operations (e.g., zooming or region enhancement) and injects intermediate visual hints into the reasoning trace. While powerful, these approaches are prone to redundant or invalid tool invocations, introducing noise that degrades performance and substantially increases inference latency. Recently, emerging researches position Latent Reasoning as a viable direction for optimizing multimodal reasoning. Unlike traditional methods that rely on explicit textual chains, latent reasoning encodes intermediate reasoning steps into compact continuous vectors. This paradigm offers compelling advantages, including higher inference efficiency, reduced annotation overhead, and the ability to learn dense, high-fidelity multimodal representations [zhang2023multimodal, DBLP:journals/corr/abs-2505-16552].
@@
To gain deeper insights into the optimization dynamics of latent reasoning, we conducted a systematic analysis of gradient flows and parametric evolution during training. This investigation reveals two meaningful observations: (1) Visual-Text Optimization Disparity: Recent studies [DBLP:journals/corr/abs-2507-03019, DBLP:conf/iclr/0002KCI25] have highlighted the phenomenon of visual attention attenuation in MLLMs as the explicit Chain-of-Thought (CoT) reasoning chain extends. We discover that a similar degradation existing in latent reasoning. As depicted in Fig. 1(a), visual tokens consistently exhibit significantly higher gradient norms and pronounced volatility compared to textual tokens. We reckon that this disparity stems from a fundamental mismatch between continuous visual features and discrete text tokens. During joint training, textual modality dominates the optimization process, causing visual representations to undergo large and erratic updates as they struggle to establish the fine-grained visual-text alignment. (2) Fixed-Depth Optimization Dilemma: Beyond modality imbalance, we observe an architecture bottleneck in how models handle samples with different complexity. Partitioning the training data into easy (consistently correct) and hard (persistently incorrect) subsets based on early-stage validation accuracy, we tracked the gradient nuclear norms across the QKV and output projection matrices O (Fig. 1(b) and Fig. 1(c)). The trends reveal a stark divergence: easy samples exhibit smooth gradient decay, indicating stable convergence into favorable loss basins. In contrast, hard samples maintain persistently high gradient volatility even in late training epochs. This phenomenon underscores the necessity of iterative refinement for complex reasoning, which is crucial for parsing compositional patterns in complex contexts while mitigating inherent visual ambiguities in images. Fixed-depth architectures, however, lack the flexibility to adapt to varying token complexities, thereby trapping hard examples in oscillatory optimization trends.
@@
Motivated by these observations, we propose a unified framework that strengthens fine-grained visual engagement and achieves token-wise depth scaling strategy to enable more precise and comprehensive contextual reasoning. Firstly, to mitigate visual optimization instability, we design a visual replay module, which dynamically replays the focused visual clues interleaved with thinking latents. This mechanism iteratively exposes and propagates key visual context across reasoning steps, fostering step-wise alignment with the target answer. Complementing this, we apply self-distillation supervision to enforce spatial coherence and preserve fine-grained visual details within the visual latents. Secondly, to address the fixed-depth optimization bottleneck, we propose a per-layer token router that dynamically allocates additional reasoning steps based on token complexity or information density. This design enables high-difficulty tokens to engage in prolonged contextual reasoning by reusing layer-wise knowledge, facilitating iterative representation refinement and adaptive prioritization of salient contextual cues. Finally, in contrast to methods [hao2024training, codi] that rely on knowledge distillation for direct latent supervision, we employ a curriculum learning strategy that progressively introduces latent tokens into the training pipeline. Extensive experiments show that our method can be effortlessly combined with various widely-used MLLM backbones to further enhance reasoning performance while maintain the satisfactory inference latency. The main contributions can be summarized as follows:
@@ 2.1 Explicit Multimodal Reasoning
Multimodal reasoning enables model to reason over information from different modalities to solve complex tasks. There are many prior works [hu2022promptcap, ssa, Zheng_NeurIPS2023, MitraCCoT, DBLP:journals/corr/abs-2601-01984, MitraCCoT, kamcot, dynfocus, super] focusing extensively on enhancing reasoning capabilities. Earlier work rely on the CoT prompting to perform explicit thinking steps in the text space before generating the final answer. However, this paradigm generally lack sufficient visual grounding capability and leads to unsatisfactory misalignment and hallucination [huang-etal-2024-visual, DBLP:journals/corr/abs-2404-18930]. To address these limitations, recent studies [Zheng_NeurIPS2023, hu2022promptcap, MitraCCoT] have explored converting visual information into textual formats prior to reasoning, leveraging external tools or specialized visual experts to generate descriptive representations that guide LLMs. For instance, hu2022promptcap pioneered the integration of visual captions, extracting semantic content as text and concatenating it with input prompts to bolster reasoning capabilities. To further enhance fine-grained reasoning, subsequent works have focused on regional understanding, aiming to improve textual expressiveness by describing specific image regions. Others [DBLP:journals/corr/abs-2601-01984, MitraCCoT, kamcot] identify entities and their relationships within images, facilitating fine-grained reasoning through explicit modeling of inter-entity connections.
@@
A line of concurrent works advocates using vision-text interleaved format during the rationale generation and reasoning process. The model draws auxiliary lines or marks based on original image to record thinking path, zoom or crop regions, or perform code editing, etc. Building on these paradigms, zhang2023multimodal first proposed decoupling rationale generation from answer generation in the Vision-Text Reasoning field. Subsequently, Shao2024VisualCA annotates key regions of the original image in intermediate steps, training models to focus on image regions relevant to the answer. While some works [gao2024interleaved, zhang2025chain] further extract key image regions progressively during reasoning, combining visual information with textual reasoning to generate the final answer. Moreover, new methods [visualsketchpad, liu2025Visual] emulate human thought by sketching images during reasoning, focusing on core concepts, structures, and relationships while ignoring redundant details. Other works [li2025imaginereasoningspacemultimodal, chern2025thinking] generate new images with auxiliary markers during reasoning, combining them with text to improve reasoning in complex scenarios. To fully shift reasoning from the linguistic domain to the visual modality, xu2025visualplanningletsthink proposes to reasoning exclusively with dynamic generated images, demonstrating substantial performance gains in visual navigation tasks.
@@
More recently, Vision-R1 [DBLP:journals/corr/abs-2503-06749] and VL-Rethinker [DBLP:journals/corr/abs-2504-08837] leverage Group Relative Policy Optimization [deepseekr1] (GRPO) to refine reasoning trajectories through rollout-based sampling and reward scoring. Complementing these policy-driven approaches, concurrent works further enhance reasoning capabilities via novel cognitive paradigms, including self-critiquing cycles [DBLP:journals/corr/abs-2405-06682, DBLP:conf/cvpr/CocchiMC0C25], iterative rethinking [lookback2025], and on-policy distillation [DBLP:journals/corr/abs-2601-18734].
2.2 Latent Reasoning
@@
Different from explicit reasoning in the discrete token space, latent reasoning refers to internal computation performed in a hidden space before answer generation. hao2024training pioneers continuous latent space reasoning by feeding the last hidden states as input embeddings for the next step without generating intermediate discrete tokens, substantially reducing reasoning latency. However, subsequent studies have indicated that such paradigm may suffer from feature homogenization without explicit supervision on intermediate latent states. To address this limitation, a series of works attempt to enhance the quality of intermediate representations using diverse strategies. cheng2024compressed introduced variable-length contemplation tokens for latent reasoning, mitigating quality degradation caused by fixed-length constraints. Similarly, shen2025codi leveraged distillation tactic to align student and teacher hidden activations along with explicit supervision, thereby constraining latent reasoning paths. Beyond these efforts, DBLP:journals/corr/abs-2509-20317 adopted step-level supervision to further stabilize the reasoning space.
@@
Recently, latent reasoning has been extended to Multimodal Large Language Models (MLLMs). Distinct from text-only LLMs, MLLMs necessitate the effective integration of visual features within the latent reasoning space. Several efforts [yang2025machine, li2025latent, pham2025multimodal, liu2025reasoningminddynamicmultimodal] have been dedicated to injecting visual cues into the latent space to facilitate visual-grounded reasoning. For instance, li2025latent emphasize image details by constructing a structured cognitive hierarchy, albeit relying on annotation-intensive multimodal reasoning data. Similarly, liu2025reasoningminddynamicmultimodal progressively select visual patches to inject into latent thinking tokens via confidence-guided policy gradient optimization. In this work, we systematically analyze gradient dynamics during latent reasoning training and reveal two critical bottlenecks towards token-wise optimization behavior.
@@ 3.1 Preliminary: Implicit and Explicit Decoding
Given a multimodal input comprising a question $\mathcal{Q}$ and an image $\mathcal{V}$, we first tokenize them into a sequence of text embeddings $\mathbf{Q}=\{\mathbf{q}_{i}\}_{i=1}^{N_{q}}$ and visual features $\mathbf{V}=\{\mathbf{v}_{i}\}_{i=1}^{N_{v}}$ via a word embedding matrix $\mathbf{E}\in\mathbb{R}^{|\mathcal{W}|\times d}$ and a pretrained visual encoder, respectively, where $|\mathcal{W}|$ denotes the vocabulary size and $d$ is the hidden dimension. Subsequently, we employ an autoregressive MLLM $\mathcal{F}_{\theta}$ to encode the concatenated input into an initial hidden state $\mathbf{H}^{(0)}\in\mathbb{R}^{P\times d}$, where $P=N_{q}+N_{v}$ represents the total number of tokens in the prefilling phase. During the decoding phase, we predetermine the number of implicit reasoning steps $T_{r}$ and explicit answer tokens $T_{a}$. The generation process consists of two following distinct stages.
Implicit Reasoning Phase. In the $t$-th implicit reasoning step, the model generates a continuous latent representation $\mathbf{z}_{t}$ conditioned on the original input sequence $\mathbf{X}=\left[\mathbf{Q}\,\|\,\mathbf{V}\right]$ and all preceding latent states. This iterative process is formulated as:
@@
where $\mathbf{Z}_{<t}=\{\mathbf{z}^{(1)},\dots,\mathbf{z}^{(t-1)}\}$ denotes the sequence of latent tokens generated in previous steps, and $\mathcal{T}(\cdot)$ extracts the final vector representation from the model output (i.e., the hidden state corresponding to the last generated token). Unlike vanilla explicit reasoning, each step yields a informative continuous latent vector rather than a discrete token. The resulting latent sequence $\mathbf{Z}_{1:T_{r}}=\{\mathbf{z}^{(t)}\}_{t=1}^{T_{r}}$ is concatenated to the context before transitioning to the explicit answer decoding phase.
Explicit Answer Decoding. In the explicit phase, the model generates discrete answer tokens by sampling from the vocabulary distribution. The conditional probability of the $t$-th answer token $a_{t}$ is given by:
$p_{\theta}(a_{t}\mid\mathbf{X},\mathbf{Z}_{1:T_{r}},a_{<t})=\mathrm{Softmax}\left(\mathbf{W}_{o}\cdot\mathbf{h}_{t}^{\text{dec}}\right),$ (2)
where $\mathbf{h}_{t}^{\text{dec}}$ is the decoder hidden state at step $t$, and $\mathbf{W}_{o}\in\mathbb{R}^{|\mathcal{W}|\times d}$ is the output projection matrix that maps hidden states to vocabulary logits.
The likelihood of the complete answer sequence $\mathbf{a}_{1:T_{a}}$ is factorized as an autoregressive product:
@@
Here, the full context input for the decoder is defined as $\mathbf{U}_{t}=\left[\mathbf{X}\,\|\,\mathbf{Z}_{1:T_{r}}\,\|\,\mathbf{a}_{<t}\right]$, where $\|$ denotes sequence concatenation along the token dimension.
Figure 2: Left Panel: Schematic illustration of our framework. Input images are encoded into visual tokens by a pretrained visual encoder and projected into a text-centric semantic space aligned with the LLM, while questions are tokenized by the text tokenizer. Our method enhances a standard latent MLLM with two synergistic components during the implicit reasoning phase: (1) Spatially-Coherent Finer Visual Replay (SCF-VR): Attention-guided selection identifies salient visual regions at each implicit step, generating fine-grained visual latents enhanced by spatially-coherent constraints. (2) Routing Depth Scaling (RDS): A lightweight learnable router adaptively allocates additional reasoning steps for high-difficulty tokens or latents. Refined hidden states and visual latents are interleaved and propagated to subsequent reasoning steps. The overall objective is jointly optimized via standard cross-entropy loss and self-distillation loss. Right Panel: Detailed architecture of the SCF-VR module and token-wise depth scaling mechanism. TokenSR denotes the token super-resolution module, and Pool is the average pooling.

### 11 May 2026 v4

@@ Abstract
Multimodal latent reasoning has emerged as a promising paradigm that replaces explicit Chain-of-Thought (CoT) decoding with implicit feature propagation, simultaneously enhancing representation informativeness and reducing inference latency. By analyzing token-level gradient dynamics during latent training, we reveal two critical observations: (1) visual tokens exhibit significantly smaller gradient norms than their textual counterparts due to inherent language bias, resulting in systematic visual under-optimization; and (2) semantically simple tokens converge rapidly, whereas complex tokens exhibit persistent gradient instability constrained by fixed architectural depths. To address these limitations, we propose a visual replay module and routing depth scaling to collaboratively enhance visual perception and refine complicated latents for deeper contextual reasoning. The former module leverages causal self-attention to estimate token saliency, reinforcing fine-grained grounding through spatially-coherent constraints. Complementarily, the latter mechanism adaptively allocates additional reasoning steps to complex tokens, enabling deeper contextual refinement. Guided by a curriculum strategy that progressively internalizes explicit CoT into compact latent representations, our framework achieves state-of-the-art performance across diverse benchmarks while delivering substantial inference speedups over explicit CoT baselines.
@@ 1 Introduction
Over the past few years, large language models (LLMs) have achieved remarkable progress in complex reasoning, propelled by scaling laws in data volume and model capacity [25, 49]. Advanced techniques such as Chain-of-Thought (CoT) prompting [80, 46] and reinforcement learning (RL) [9, 73, 12] for trajectory optimization have proven highly effective in text-only domains. Extending these capabilities to the multimodal realm has thus become a pivotal research direction. Current approaches primarily follow three paradigms. First, Text-based Reasoning [80, 39, 26] generates explicit multi-step textual chains before producing an answer. However, these methods typically rely on static visual inputs, and recent studies [69, 74] indicate that visual grounding deteriorates significantly over extended reasoning chains. Second, Tool-augmented Reasoning manipulates visual inputs through external operations (e.g., zooming or region enhancement) and injects intermediate visual hints into the reasoning trace. While powerful, these approaches are prone to redundant or invalid tool invocations, introducing noise that degrades performance and substantially increases inference latency. Recently, emerging researches position Latent Reasoning as a viable direction for optimizing multimodal reasoning. Unlike traditional methods that rely on explicit textual chains, latent reasoning encodes intermediate reasoning steps into compact continuous vectors. This paradigm offers compelling advantages, including higher inference efficiency, reduced annotation overhead, and the ability to learn dense, high-fidelity multimodal representations [80, 51].
@@
Figure 1: Panel (a) depicts the token-wise Frobenius norm of gradients across different layers throughout the entire training process, revealing a clear gradient imbalance between visual tokens and text tokens/latent token. Panels (b) and (c) further reveal distinct evolution patterns of the Nuclear norm in the QKVO projection matrix across layers and training epochs: gradients for easy samples decay rapidly and converge smoothly, whereas hard samples maintain elevated gradient norms with persistent oscillations. Here, $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$ denote the low-rank matrix in LoRA [19].
@@
To gain deeper insights into the optimization dynamics of latent reasoning, we conducted a systematic analysis of gradient flows and parametric evolution during training. This investigation reveals two meaningful observations: (1) Visual-Text Optimization Disparity: Recent studies [68, 77] have highlighted the phenomenon of visual attention attenuation in MLLMs as the explicit Chain-of-Thought (CoT) reasoning chain extends. We discover that a similar degradation existing in latent reasoning. As depicted in Fig. 11(a), the gradient updates for visual tokens (regions marked by dashed lines) are remarkably small, while the overall optimization is predominantly governed by gradients from text tokens and latent token. This suggests that the model relies primarily on textual information and latent representations to generate final answers during training, revealing a pronounced gradient imbalance phenomenon. (2) Fixed-Depth Optimization Dilemma: Beyond modality imbalance, we observe an architecture bottleneck in how models handle samples with different complexity. Partitioning the training data into easy (consistently correct) and hard (persistently incorrect) subsets based on early-stage validation accuracy, we tracked the gradient nuclear norms across the QKV and output projection matrices O (Fig. 11(b) and Fig. 11(c)). The trends reveal a stark divergence: easy samples exhibit smooth gradient decay, indicating stable convergence into favorable loss basins. In contrast, hard samples maintain persistently high gradient volatility even in late training epochs. This phenomenon underscores the necessity of iterative refinement for complex reasoning, which is crucial for parsing compositional patterns in complex contexts while mitigating inherent visual ambiguities in images. Fixed-depth architectures, however, lack the flexibility to adapt to varying token complexities, thereby trapping hard examples in oscillatory optimization trends.
@@
Motivated by these observations, we propose a unified framework that strengthens fine-grained visual engagement and achieves token-wise depth scaling strategy to enable more precise and comprehensive contextual reasoning. Firstly, to mitigate visual optimization instability, we design a visual replay module, which dynamically replays the focused visual clues interleaved with thinking latents. This mechanism iteratively exposes and propagates key visual context across reasoning steps, fostering step-wise alignment with the target answer. Complementing this, we apply self-distillation supervision to enforce spatial coherence and preserve fine-grained visual details within the visual latents. Secondly, to address the fixed-depth optimization bottleneck, we propose a per-layer token router that dynamically allocates additional reasoning steps based on token complexity or information density. This design enables high-difficulty tokens to engage in prolonged contextual reasoning by reusing layer-wise knowledge, facilitating iterative representation refinement and adaptive prioritization of salient contextual cues. Finally, in contrast to methods [18, 47] that rely on knowledge distillation for direct latent supervision, we employ a curriculum learning strategy that progressively introduces latent tokens into the training pipeline. Extensive experiments show that our method can be effortlessly combined with various widely-used MLLM backbones to further enhance reasoning performance while maintain the satisfactory inference latency. The main contributions can be summarized as follows:
@@ 2 Related Work
Explicit Multimodal Reasoning. Multimodal reasoning enables model to reason over information from different modalities to solve complex tasks. There are many prior works [75, 20, 16, 82, 39, 38, 39, 40, 15, 17, 76, 35, 70, 71] focusing extensively on enhancing reasoning capabilities. Earlier work rely on the CoT prompting to perform explicit thinking steps in the text space before generating the final answer. However, this paradigm generally lack sufficient visual grounding capability and leads to unsatisfactory misalignment and hallucination [22, 3]. To address these limitations, recent studies [82, 20, 39] have explored converting visual information into textual formats prior to reasoning, leveraging external tools or specialized visual experts to generate descriptive representations that guide LLMs. For instance, Hu et al. [20] pioneered the integration of visual captions, extracting semantic content as text and concatenating it with input prompts to bolster reasoning capabilities. To further enhance fine-grained reasoning, subsequent works have focused on regional understanding, aiming to improve textual expressiveness by describing specific image regions. Others [38, 39, 40] identified entities and their relationships within images, facilitating fine-grained reasoning through explicit modeling of inter-entity connections.
@@
A line of concurrent works advocates using vision-text interleaved format during the rationale generation and reasoning process. The model draws auxiliary lines or marks based on original image to record thinking path, zoom or crop regions, or perform code editing, etc. Building on these paradigms, Zhang et al. [80] first proposed decoupling rationale generation from answer generation in the Vision-Text Reasoning field. Subsequently, Shao et al. [45] annotated key regions of the original image in intermediate steps, training models to focus on image regions relevant to the answer. While some works [13, 79] further extract key image regions progressively during reasoning, combining visual information with textual reasoning to generate the final answer. Moreover, new methods [21, 34] emulated human thought by sketching images during reasoning, focusing on core concepts, structures, and relationships while ignoring redundant details. Other works [31, 7] generated new images with auxiliary markers during reasoning, combining them with text to improve reasoning in complex scenarios. To fully shift reasoning from the linguistic domain to the visual modality, Xu et al. [67] proposes to reasoning exclusively with dynamic generated images, demonstrating substantial performance gains in visual navigation tasks.
@@
More recently, Vision-R1 [23] and VL-Rethinker [54] leveraged Group Relative Policy Optimization [9] (GRPO) to refine reasoning trajectories through rollout-based sampling and reward scoring. Complementing these policy-driven approaches, concurrent works further enhance reasoning capabilities via novel cognitive paradigms, including self-critiquing cycles [44, 8], iterative rethinking [69], and on-policy distillation [81].
@@
Latent Reasoning. Different from explicit reasoning in the discrete token space, latent reasoning refers to internal computation performed in a hidden space before answer generation. Hao et al. [18] pioneers continuous latent space reasoning by feeding the last hidden states as input embeddings for the next step without generating intermediate discrete tokens, substantially reducing reasoning latency. However, subsequent studies have indicated that such paradigm may suffer from feature homogenization without explicit supervision on intermediate latent states. To address this limitation, a series of works attempt to enhance the quality of intermediate representations using diverse strategies. Cheng and Van Durme [6] introduced variable-length contemplation tokens for latent reasoning, mitigating quality degradation caused by fixed-length constraints. Similarly, Shen et al. [48] leveraged distillation tactic to align student and teacher hidden activations along with explicit supervision, thereby constraining latent reasoning paths. Beyond these efforts, Wei et al. [65] adopted step-level supervision to further stabilize the reasoning space.
@@
Recently, latent reasoning has been extended to Multimodal Large Language Models (MLLMs). Distinct from text-only LLMs, MLLMs necessitate the effective integration of visual features within the latent reasoning space. Several efforts [72, 28, 43, 33, 62] have been dedicated to injecting visual cues into the latent space to facilitate visual-grounded reasoning. For instance, Wang et al. [62] emphasize image details by constructing a structured cognitive hierarchy, albeit relying on annotation-intensive multimodal reasoning data. Similarly, Liu et al. [33] progressively select visual patches to inject into latent thinking tokens via confidence-guided policy gradient optimization. In this work, we systematically analyze gradient dynamics during latent reasoning training and reveal two critical bottlenecks towards token-wise optimization behavior.
@@ 3.1 Preliminary: Implicit and Explicit Decoding
Given a multimodal input comprising a question $\mathcal{Q}$ and an image $\mathcal{V}$, we tokenize them into a text embedding sequence $\mathbf{Q}=\{\mathbf{q}_{i}\}_{i=1}^{N_{q}}$ and visual features $\mathbf{V}=\{\mathbf{v}_{i}\}_{i=1}^{N_{v}}$ using a word embedding matrix $\mathbf{E}\in\mathbb{R}^{|\mathcal{W}|\times d}$ and a pretrained visual encoder, respectively. We concatenate $\mathbf{X}=[\mathbf{Q}\|\mathbf{V}]$ and use an autoregressive MLLM $\mathcal{F}_{\theta}$ to obtain the prefilling hidden states $\mathbf{H}^{(0)}\in\mathbb{R}^{P\times d}$, where $P=N_{q}+N_{v}$. During decoding, we fix $T_{r}$ implicit reasoning steps and $T_{a}$ explicit answer tokens. The generation process has two stages.
Implicit Reasoning Phase. At step $t$, the model produces a continuous latent vector $\mathbf{z}^{(t)}$ conditioned on $\mathbf{X}$ and previous latent states:
@@
where $\mathbf{Z}_{<t}=\{\mathbf{z}^{(1)},\dots,\mathbf{z}^{(t-1)}\}$ and $\mathcal{T}(\cdot)$ extracts the hidden state of the last token. The latent sequence $\mathbf{Z}_{1:T_{r}}$ is then concatenated to the context for answer decoding.
Explicit Answer Decoding. The answer tokens $\mathbf{a}_{1:T_{a}}$ are generated autoregressively. For $t$-th answer token $a_{t}$:
$p_{\theta}(a_{t}\mid\mathbf{X},\mathbf{Z}_{1:T_{r}},a_{<t})=\mathrm{Softmax}\!\left(\mathbf{W}_{o}\mathbf{h}_{t}^{\text{dec}}\right),$ (2)
where $\mathbf{W}_{o}\in\mathbb{R}^{|\mathcal{W}|\times d}$ is the output projection matrix and $\mathbf{h}_{t}^{\text{dec}}$ is the decoder hidden state at step $t$. The joint likelihood factorizes as
@@
At each decoding step, the decoder input is $\mathbf{U}_{t}=[\mathbf{X}\|\mathbf{Z}_{1:T_{r}}\|a_{<t}]$, with $\|$ denoting concatenation along the token dimension.
Figure 2: Left Panel: Schematic illustration of our framework. Input images are encoded into visual tokens by a pretrained visual encoder and projected into a text-centric semantic space aligned with the LLM, while questions are tokenized by the text tokenizer. Our method enhances a standard latent MLLM with two synergistic components during the implicit reasoning phase: Spatially-Coherent Finer Visual Replay (SCF-VR) and Routing Depth Scaling (RDS). The overall objective is jointly optimized via standard cross-entropy loss and self-distillation loss. Right Panel: Detailed architecture of the SCF-VR module and token-wise depth scaling mechanism. TokenSR denotes the token super-resolution module, and Pool is the average pooling.